40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 23:14:24 2018, aaron, Update, OMC, OMC scanning/aligning script 
    Reply  Tue Nov 20 20:33:11 2018, aaron, Update, OMC, OMC scanning/aligning script pathwaysketch.pdfIMG_0094.JPGIMG_0097.JPG
       Reply  Mon Nov 26 15:43:16 2018, aaron, Update, OMC, OMC scanning/aligning script c1omc.png
          Reply  Thu Dec 6 11:16:28 2018, aaron, Update, OMC, OMC channels 
             Reply  Mon Dec 10 12:29:05 2018, aaron, Update, OMC, OMC channels 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg545E5512-D003-408B-9F00-55F985966A16.jpegDFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
                Reply  Mon Dec 10 19:47:06 2018, aaron, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:42:44 2018, Koji, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:48:04 2018, aaron, Update, OMC, OMC channels 
                      Reply  Thu Dec 13 22:24:21 2018, aaron, Update, OMC, OMC channels 
          Reply  Mon Dec 10 12:11:28 2018, aaron, Update, OMC, Aligning the OMC B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
             Reply  Tue Dec 11 14:24:18 2018, aaron, Update, OMC, Aligning the OMC 
                Reply  Tue Dec 11 22:50:07 2018, aaron, Update, OMC, Aligning the OMC 6x
                   Reply  Thu Dec 13 22:36:42 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:02:24 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:06:12 2018, aaron, Update, OMC, Aligning the OMC 
                         Reply  Mon Dec 17 20:45:40 2018, aaron, Update, OMC, Aligning the OMC IMG_5922.JPGIMG_5921.JPG
                            Reply  Tue Dec 18 18:13:32 2018, aaron, Update, OMC, OMC L HV Piezo driver tests (again) IMG_4576_copy.jpg
                               Reply  Wed Dec 19 00:12:46 2018, gautam, Update, OMC, 40m OMC DCC node 
Message ID: 14357     Entry time: Fri Dec 14 13:02:24 2018     In reply to: 14355
Author: aaron 
Type: Update 
Category: OMC 
Subject: Aligning the OMC 

I replaced the 2'' AUX-AS combining BS with a 2'' HR mirror for 1064. I aligned the AUX beam from the new HR mirror into the next iris, so AUX passes through irises both before and after the new optic. Now, AS does not go out to the AS PDs.

I mounted the old BS on the SP table in a random orientation.

I also dumped the beam transmitted through one of the AUX steering mirrors before the new HR mirror.

ELOG V3.1.3-